solve each system of equation to identify the (xy) solution set
y=5x-11
y=2x+10

Answer by venugopalramana(3286)

solve each system of equation to identify the (xy) solution set
y=5x-11................1
y=2x+10...............................2
substitute for y in eqn.2 from eqn.2
5x-11=2x+10
5x-2x=10+11
3x=21
x=7....hence
y=2x+10=2*7+10=14+10=24
(7,24) is the solution set.